OpenStudy (anonymous):

find the value of theta, sec theta=2

OpenStudy (unklerhaukus):

\[\sec=\frac1\cos\]

OpenStudy (anonymous):

secθ = 1/cosθ = 2 cosθ = 1/2 According to unit circle, cos(±60°)=1/2, so θ = 60°, 300